How much heat is released when an acetylene (C2H2) torch burns 250 g of fuel? ∆H= -1300 kJ?


1
Expert's answer
2021-02-11T04:49:16-0500

C2H2 + 3/2O2 = CO2 + H2O + 1300 kJ.

1300 x 250/26 = 12 500 kJ.